OpenStudy (anonymous):

Rewrite using scientific notation: 338 A: 3.38 × 10^1 B:3.38 × 10^2 C:3.38 × 10^3 D:3.38 × 10^4

OpenStudy (anonymous):

See, when we multiply and divide some number by same number, then its value does not change.

OpenStudy (lxelle):

3.38 x 10^2 = 338 by moving the decimal to the right twice.

OpenStudy (anonymous):

For example : \[5 \] Multiply and divide it by 10: \[5 \times \frac{10}{10}\] = \(5\) only

OpenStudy (anonymous):

Coming to your problem, you have \(338\).. You want to shift decimal point from right to left, so we can divide and multiply by \(100\) here.

OpenStudy (anonymous):

\[338 \times \frac{100}{100} = \frac{338}{100} \times 100\]

OpenStudy (anonymous):

We can use \(100\) for division there;\[\implies 3.38 \times 100 = \color{green}{3.38 \times 10^2}\]